# Training in reverse: How iteration order influences convergence and stability in deep learning

### v1

@@ Abstract
Despite exceptional achievements, training neural networks remains computationally expensive and is often plagued by instabilities that can degrade convergence. While learning rate schedules can help mitigate these issues, finding optimal schedules is time-consuming and resource-intensive. This work explores theoretical issues concerning training stability in the constant-learning-rate (i.e., without schedule) and small-batch-size regime. Surprisingly, we show that the order of gradient updates affects stability and convergence in gradient-based optimizers. We illustrate this new line of thinking using backward-SGD, which processes batch gradient updates like SGD but in reverse order. Our theoretical analysis shows that in contractive regions (e.g., around minima) backward-SGD converges to a point while the standard forward-SGD generally only converges to a distribution. This leads to improved stability and convergence which we demonstrate experimentally. While full backward-SGD is computationally intensive in practice, it highlights opportunities to exploit reverse training dynamics (or more generally alternate iteration orders) to improve training. To our knowledge, this represents a new and unexplored avenue in deep learning optimization.
**footnotetext: Equal contribution; corresponding author dherin@google.com§§footnotetext: Supported in part by the Swiss NSF grants 200020-200400, 200021-212864 and the Swedish Research Council grant 104651320.
@@ 1 Introduction
In recent years, neural networks have achieved remarkable success across diverse domains from text generation [1, 2, 3] and image creation [4, 5, 6] to applications in protein folding [7, 8] and material discovery [9]. However, their training remains challenging and computationally expensive. One of the reasons for this is due to training instabilities which often occur [10, 11, 12, 13] and which produces hard to interpret loss curves, wasted computation time, and potentially failed experiments. One way to view this challenge is as a trade-off between stability and performance: hyperparameter settings that often yield better test performance, such as higher learning rates and smaller batch sizes, tend to exacerbate these instabilities.
@@
As a basic example, consider the training batch size. For many common optimizers, smaller batch sizes often lead to improved test performance. In fact, recent research has shown that small batches induce a form of implicit regularization [14, 15, 16, 17, 18] which benefits generalization. On the other hand, the greater variability of small batches tends to exacerbate oscillations of the training loss, prolonging time to convergence. In this work, we demonstrate that these instability and convergence issues associated with small batch sizes can be mitigated by applying the optimizer in reverse; i.e., iterating through batches in a backward order, which we now explain.
@@
Namely, standard training algorithms, such as stochastic gradient descent (SGD), Adam, and other gradient-based optimizers, are iterative processes. At each step, these algorithms update the network parameters $\theta$ using a randomly sampled data batch $B_{i}$. This update can be formalized as a transformation, $\theta^{\prime}=T_{i}(\theta)$, where $\theta^{\prime}$ represents the new parameter value, aiming to reduce the training loss on batch $B_{i}$. Because at each step $i$ the batch $B_{i}$ is randomly sampled, the update operator $T_{i}$ can be formalized as a random operator. For instance, in the case of SGD, the random operator is defined by the update rule $T_{i}(\theta)=\theta-h\nabla L_{B_{i}}(\theta)$, where $L_{B_{i}}$ is the loss function evaluated on batch $B_{i}$ and $h$ is the learning rate. The sequence of parameter updates generated by these iterations, starting from an initial parameter value $\theta$, defines the standard learning trajectory, which we will refer to as the forward trajectory:
$\theta,\quad T_{1}(\theta),\quad T_{2}T_{1}(\theta),\quad\dots,\quad T_{n}T_{n-1}\cdots T_{1}(\theta).$
@@
Moderate or large learning rates and small batches, when used with the standard forward trajectory for training, tend to destabilize common gradient-based optimizers, causing convergence and stability issues. Consequently, learning rate schedules have become essential as an attempt to encourage proper convergence of the loss during training. Our key contribution, which we support with both theoretical analysis and empirical results (see Figure 2 and Appendix A), is in demonstrating that reversing the iteration order–what we call the backward trajectory–leads to significantly more stable convergence. Specifically, the backward trajectory consists of iterates generated by applying the training batches in reverse order; i.e.,
$\theta,\quad T_{1}(\theta),\quad T_{1}T_{2}(\theta),\quad\dots,\quad T_{1}T_{2}\cdots T_{n}(\theta).$
For the sake of clarity, let us work out the first two iterates both for forward and backward. For instance for SGD, if we receive first the batch $B_{1}$ and then the batch $B_{2}$ in sequence in the training loop, then at the first step both the backward and forward iterates coincides with $T_{1}(\theta)=\theta-h\nabla L_{B_{1}}(\theta)$, where $\theta$ in the randomly initialized parameter value. However at the second step the forward iterate becomes
@@
A naive (and computation intensive) implementation of the backward optimization is depicted in Figure 1 where all the batches received so far at each steps are re-processed from scratch in the reverse order in which they were received from the initialization point.
@@
The main contributions of this paper are to show theoretically (Theorem 2.2 and Theorem 2.6) and experimentally (Figure 2 and Appendix A) that backward optimization has better convergence and stability properties than the standard forward optimization. As already known (see related work in Section 1.1), forward trajectories do not generally converge toward points but rather toward a probability distribution of the iterates (in the fixed learning-rate regime).
We argue that the advantage of backward trajectories comes from their convergence toward actual points (see Theorem 2.2) sampled from the forward distribution (see Theorem 2.6). We prove this using a generalization of the Banach fixed point theorem when the random maps $T_{i}$ become contractions. Note that backward optimization can theoretically be applied to any gradient-based optimizer (not only to SGD) to improve stability and convergence, since our theoretical statements hold at the level of the $T_{i}$’s, no matter what their actual forms are. To our knowledge, these concepts are new in deep learning optimization. The main goal of this paper is to expose how this phenomenon manifests in deep learning. We defer engineering applications leveraging this phenomenon (like efficient implementations of the backward SGD) to future work, while outlining a few potential directions at the paper’s conclusion.
@@ 1.1 Related work
A number of works prove under different assumptions that the (forward) iterates of SGD with constant learning rate do not converge toward points but rather toward a stationary probability distribution: [19] and [20] show this for the strongly convex case; [21] treats the non-convex case with separable loss; [22] focus on losses coming from exponential family models; [23] quantify the rate of convergence of SGD to its stationary distribution in non-convex optimization problems; see also [24] and [25]. In [26], the authors leverage the fact that SGD with fixed large learning rate oscillates between different solutions in order to create a cheap average of models by saving the explored parameters along the way. For convergence to a particular solution, the forward order for SGD needs an extra decaying learning rate schedule as shown in [27] or [28].
@@
From our point of view, one important feature that leads to convergence for the backward trajectory is the contraction property of the random operators. This notion (which we believe is under-exploited in deep learning) has surfaced in different contexts in deep learning: see [29], [30], and [31].
@@
In many Markov Chain Monte Carlo (MCMC) algorithms the goal is to sample from a distribution $\mu$. The idea is to construct a Markov chain with stationary distribution $\mu$ and then run the chain for a long time to get samples from $\mu$. The Propp-Wilson algorithm uses a form of backward iterations to accelerate the convergence toward samples from the distribution (see [32]). More generally, the idea of backward dynamics is hidden in many constructions in Markov chain theory when the Markov chain is given by iterations of random operators as illustrated in [33]. In particular, they prove a general result concerning the convergence in distribution of these types of iterated Markov chains using the backward dynamics.
@@
Already instability issues appear in the full-batch regime, and a number of theoretical works have studied it under the heading edge of stability [23, 34, 35]. Other works have also studied stability in the large batch regime after a batch size saturation takes place using the implicit conditioning ratio [36, 37]. In this work we focus on the stochastic or small batch setting. Our findings do not really matter for the full-batch setting since backward and forward iterates then coincide. In the context of physics-informed neural networks, it has been observed that the gradient field is stiff, producing instabilities in learning trajectories [38]. To remedy this [39] propose a backward Euler scheme to stabilize training in this context. However the backward Euler method is an implicit Runge-Kutta method, which is different from using iteration backward.
@@
Curriculum learning [40] leverages the impact of sample order for generalization [41] by organizing training examples in a meaningful sequence, typically starting with simpler examples and gradually introducing harder ones, thereby optimizing the learning process and improving model performance. The backward optimization can be viewed as an automated form of curriculum, mitigating the forgetting of previous examples as new examples are added; a phenomenon which is also related to catastrophic forgetting and the stability gap in continual learning (see [42] for instance).
@@ 2 A backward contraction principle
The contraction mapping principle, also called the Banach fixed point theorem, is a cornerstone result in mathematics and science, in particular for finding solutions of equations (Newton’s method) or of differential equations (Picard’s method of successive approximation). It is also behind Google’s PageRank algorithm [43]. It concerns the existence of a fixed point of maps which are uniform contractions of a complete metric space §§§Recall that a complete metric space $(\Omega,d)$ is a set $\Omega$, equipped with a distance metric $d(\theta_{1},\theta_{2})$ for which all Cauchy sequences (i.e., sequences $\theta_{n}$ such that $d(\theta_{n},\theta_{m})\rightarrow 0$ as $n,m\rightarrow 0$) converge to points in the space; typically $\mathbb{R}^{d}$ with the Euclidean distance is a complete metric space.. In detail, let $(\Omega,d)$ be a complete metric space and $T:\Omega\rightarrow\Omega$ be a continuous map. If there exists $10\leq k<$ 1 such that
@@
for any $\theta\in\Omega$ as $n\rightarrow\infty$. For example, in PageRank [43], one iterates the PageRank matrix from 50 to 100 times and the result is a very good approximation of the PageRank vector $\theta_{0}$. Another application is the standard proof of convergence for (full-batch) gradient descent around a minimum, which can be interpreted as an application of the Banach fixed point theorem as detailed in the next example:
@@
(Full-batch gradient descent convergence.) In this case the operator is $T(\theta)=\theta-h\nabla L(\theta)$ for a loss function $L$. The idea of the proof is to choose the learning rate $h$ small enough so that $T$ becomes a contraction. More precisely, around a minimum $\theta^{*}$ the operator $T$ can be approximated using a first-order expansion of the gradient around the minimum as $T(\theta)=\theta^{*}+(1-hH)(\theta-\theta^{*}),$ where $H=\nabla^{2}L(\theta^{*})$. Now we have that
$\|T(\theta_{1})-T(\theta_{2})\|\leq\|1-hH\|_{\textrm{op}}\|\theta_{1}-\theta_{2}\|,$ (1)
where $\|1-hH\|_{\textrm{op}}$ is the operator norm of $1-hH$, that is the operator maximum eigenvalue: $\max_{i}|1-h\lambda_{i}|$ (here the $\lambda_{i}$’s are eigenvalues of $H$). It is easy to verify that $\|1-hH\|_{\textrm{op}}<1$ if and only if the learning rate is strictly smaller than $2/\lambda_{\textrm{max}}$ where $\lambda_{\textrm{max}}$ is the largest eigenvalue of $H$. Convergence for that setting follows from the Banach fixed point theorem.
The generalization of this convergence argument to Stochastic Gradient Descent (SGD) is actually problematic. The main reason is that now we do not have a single operator but a sequence of them $T_{1},T_{2},\dots$, each computing the loss gradient on a different batch of data. Even if each of the operators is a contraction it turns out that usual forward iterations will not converge to a point in general as illustrated in Example 2.4. Theorem 2.2, however, establishes the convergence of the backward iterations.
@@
Condition 2 expresses that the distance $d(\theta,T_{i}(\theta))$ is uniformly bounded by a constant $D$ for all $T_{i}^{\prime}s$ for a point $\theta$. Let us show that if this happens for a single point $\theta$, this happens for all points, provided we change the constant $D$. To see this, take another point $\tilde{\theta}$. We will compute another constant $\tilde{D}$ such that $d(\tilde{\theta},T_{i}(\tilde{\theta}))<\tilde{D}$. Namely, using the triangle inequality and the fact that $T_{i}$ are uniform contractions, we obtain that
$d(\tilde{\theta},T_{i}(\tilde{\theta}))$$\leq$ $d(\tilde{\theta},\theta)+d(\theta,T_{i}(\theta))+d(T_{i}(\theta),T_{i}(\tilde{\theta}))$
@@
At this point, we can use the condition in (2) for the first term $d(\theta,T_{n+1}(\theta))<D$ and in conjunction with the contraction property for the subsequent terms in the sum, yielding:
@@
Let us verify that the condition in 2 is satisfied for SGD with a loss of form $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}L_{i}(\theta)$, where $L_{i}$ is the loss computed on batch $B_{i}$. We further assume that the $L_{i}(\theta)$ are positive and that there exists at least one point $\theta^{*}$ where the full loss vanishes (which implies that $L_{i}(\theta^{*})=0$ for all $i$’s). Both assumptions are completely reasonable for over-parameterized models like deep neural networks. For SGD, the operators are $T_{i}(\theta)=\theta-\nabla L_{i}(\theta)$. Now it is easy to see that condition in 2 is satisfied: $d(\theta^{*},T_{i}(\theta^{*}))=\|\theta^{*}-\theta^{*}-h\nabla L_{i}(\theta^{*})\|=0,$ since the global zero $\theta^{*}$ is a critical point for all the $L_{i}$’s.
Now here is a major point: The contraction mapping principle fails with the forward order. As the next example demonstrates, it is not enough for the operator $T_{i}$’s to be uniform contractions for the forward sequence of iterates to converge, while this is always true for the backward sequence because of Theorem 2.2.
@@
(Forward iterations counter-example.) Here is an extreme example illustrating the convergence failure for the forward sequence of iterates (even when the maps are uniform contractions), while the backward sequence converges to a single point under the conditions of Theorem 2.2. Consider the two constant maps $S(\theta)=x_{0}$, and $U(\theta)=y_{0}$, which are contractions with $k=0$. Assume that $x_{0}\neq y_{0}$, now also assume that each $T_{i}$ is either equal to $S$ or $U$ with equal probability to be selected. Then, independently of $\theta$, the forward iterates
@@
will jump between $x_{0}$ and $y_{0}$, according to whether $T_{n}$ is $S$ or $U$ at that particular $n$. Thus no convergence to a single point is possible, although the sequence converges to a probability uniformly distributed in the two outcomes, since for each forward iterate either outcome has probability $1/2$. On the contrary, the backward trajectory

### v2

# How iteration composition influences convergence and stability in deep learning

@@ Abstract
Despite exceptional achievements, training neural networks remains computationally expensive and is often plagued by instabilities that can degrade convergence. While learning rate schedules can help mitigate these issues, finding optimal schedules is time-consuming and resource-intensive. This work explores theoretical issues concerning training stability in the constant-learning-rate (i.e., without schedule) and small-batch-size regime. Surprisingly, we show that the composition order of gradient updates affects stability and convergence in gradient-based optimizers. We illustrate this new line of thinking using backward-SGD, which produces parameter iterates at each step by reverting the usual forward composition order of batch gradients. Our theoretical analysis shows that in contractive regions (e.g., around minima) backward-SGD converges to a point while the standard forward-SGD generally only converges to a distribution. This leads to improved stability and convergence which we demonstrate experimentally. While full backward-SGD is computationally intensive in practice, it highlights that the extra freedom of modifying the usual iteration composition by reusing creatively previous batches at each optimization step may have important beneficial effects in improving training. Our experiments provide a proof of concept supporting this phenomenon. To our knowledge, this represents a new and unexplored avenue in deep learning optimization.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
In recent years, neural networks have achieved remarkable success across diverse domains from text generation Gemini (2024; 2023); Brown et al. (2020) and image creation Ramesh et al. (2022; 2021); Saharia et al. (2022) to applications in protein folding Jumper et al. (2021a; b) and material discovery Merchant et al. (2023). However, their training remains challenging and computationally expensive. One of the reasons for this is due to training instabilities which often occur Morchdi et al. (2023); Li et al. (2019); Cohen et al. (2021); Chen et al. (2018) and which produces hard to interpret loss curves, wasted computation time, and potentially failed experiments. One way to view this challenge is as a trade-off between stability and performance: hyperparameter settings that often yield better test performance, such as higher learning rates and smaller batch sizes, tend to exacerbate these instabilities.
@@
As a basic example, consider the training batch size. For many common optimizers, smaller batch sizes often lead to improved test performance. In fact, recent research has shown that small batches induce a form of implicit regularization Novack et al. (2023); Smith et al. (2021); Dherin et al. (2022); Keskar et al. (2017); Ali et al. (2020) which benefits generalization. On the other hand, the greater variability of small batches tends to exacerbate oscillations of the training loss, prolonging time to convergence. In this work, we demonstrate that these instability and convergence issues associated with small batch sizes can be mitigated without loss of generalization power by reversing the composition order used to produce each iterate. The general line of thinking behind our approach consists of leveraging not only the current batch, but also the previous batches to produce the current parameter iterate. The creative reuse of previous batches at a given step has been already explored in Choi et al. (2019) for instance in order to speed up training by composing previous batch gradient updates to the current iterate while waiting for the data pipeline to deliver a new batch in case of IO bound situations. In this work, we reverse the composition order of all the batches received so far from initialization and show theoretically and practically that the sequence of iterates produced this way enjoys better convergence and stability properties than the usual composition order of the batch gradient updates. We now describe below this reverse composition procedure in details.
@@
Namely, standard training algorithms, such as stochastic gradient descent (SGD), Adam, and other gradient-based optimizers, are iterative processes. At each step, these algorithms update the network parameters $\theta$ using a randomly sampled data batch $B_{i}$. This update can be formalized as a transformation, $\theta^{\prime}=T_{i}(\theta)$, where $\theta^{\prime}$ represents the new parameter value. Because at each step $i$ the batch $B_{i}$ is randomly sampled, the update operator $T_{i}$ can be formalized as a random operator. For instance, in the case of SGD, the random operator is defined by the update rule $T_{i}(\theta)=\theta-h\nabla L_{B_{i}}(\theta)$, where $L_{B_{i}}$ is the loss function evaluated on batch $B_{i}$ and $h$ is the learning rate. The sequence of parameter updates generated by these iterations, starting from an initial parameter value $\theta$, defines the standard learning trajectory, which we will refer to as the forward trajectory:
$\theta_{0}=\theta,\quad\theta_{1}=T_{1}(\theta),\quad\theta_{2}=T_{2}T_{1}(\theta),\quad\dots,\quad\theta_{n}=T_{n}T_{n-1}\cdots T_{1}(\theta).$
@@
Moderate or large learning rates and small batches, when used with the standard forward trajectory for training, tend to destabilize common gradient-based optimizers, causing convergence and stability issues. Consequently, learning rate schedules have become essential as an attempt to encourage proper convergence of the loss during training. Our key contribution, which we support with both theoretical analysis and empirical results (see Figures˜2 and A), is in demonstrating that reversing the batch composition order to produce an iterate at a given step–what we call the backward trajectory–leads to significantly more stable convergence. Specifically, the backward trajectory consists of iterates generated by composing the training batches in reverse order; i.e.,
$\theta_{0}=\theta,\quad\theta_{1}=T_{1}(\theta),\quad\theta_{2}=T_{1}T_{2}(\theta),\quad\dots,\quad\theta_{n}=T_{1}T_{2}\cdots T_{n}(\theta).$
For the sake of clarity, let us work out the first two iterates both for forward and backward. For instance for SGD, if we receive first the batch $B_{1}$ and then the batch $B_{2}$ in sequence in the training loop, then at the first step both the backward and forward iterates coincides with $T_{1}(\theta)=\theta-h\nabla L_{B_{1}}(\theta)$, where $\theta$ is the randomly initialized parameter value. However at the second step the forward iterate becomes
@@
A naive (and computationally intensive) implementation of the backward optimization is depicted in Figure˜1 where all the batches received so far at each step are re-processed from scratch in the reverse order in which they were received from the initialization point.
@@
The main contributions of this paper are to show theoretically (Theorems˜2.2 and 2.6) and experimentally (Figures˜2 and A) that backward optimization has better convergence and stability properties than the standard forward optimization. As already known (see related work in Section˜1.1), forward trajectories do not generally converge toward points but rather toward a probability distribution of the iterates (in the fixed learning-rate regime).
We argue that the advantage of backward trajectories comes from their convergence toward actual points (see Theorem˜2.2) sampled from the forward distribution (see Theorem˜2.6). We prove this using a generalization of the Banach fixed point theorem when the random maps $T_{i}$ become contractions. Note that backward optimization can theoretically be applied to any gradient-based optimizer (not only to SGD) to improve stability and convergence, since our theoretical statements hold at the level of the $T_{i}$’s, no matter what their actual forms are. To our knowledge, these concepts are new in deep learning optimization. The main goal of this paper is to expose how this phenomenon manifests in deep learning. We defer engineering applications leveraging this phenomenon (like efficient implementations of the backward SGD) to future work, while outlining a few potential directions at the paper’s conclusion.
Overall, this work should be understood as a proof of concept showing that reversing the iteration composition order can significantly improve stability and convergence in stochastic optimization.
@@ 1.1 Related work
A number of works prove under different assumptions that the (forward) iterates of SGD with constant learning rate do not converge toward points but rather toward a stationary probability distribution: Merad and Gaïffas (2023) and Dieuleveut et al. (2020) show this for the strongly convex case; Shirokoff and Zaleski (2024) treats the non-convex case with separable loss; Babichev and Bach (2018) focus on losses coming from exponential family models; Cheng et al. (2020) quantify the rate of convergence of SGD to its stationary distribution in non-convex optimization problems; see also Dieuleveut and Bach (2016) and Meyn and Tweedie (1993). In Huang et al. (2017), the authors leverage the fact that SGD with fixed large learning rate oscillates between different solutions in order to create a cheap average of models by saving the explored parameters along the way. For convergence to a particular solution, the forward order for SGD needs an extra decaying learning rate schedule as shown in Robbins and Monro (1951) or Mertikopoulos et al. (2020).
@@
From our point of view, one important feature that leads to convergence for the backward trajectory is the contraction property of the random operators. This notion (which we believe is under-exploited in deep learning) has surfaced in different contexts in deep learning: see Bergomi et al. (2019)Qian and Wegman (2019), and Avelin and Karlsson (2022).
@@
In many Markov Chain Monte Carlo (MCMC) algorithms the goal is to sample from a distribution $\mu$. The idea is to construct a Markov chain with stationary distribution $\mu$ and then run the chain for a long time to get samples from $\mu$. The Propp-Wilson algorithm uses a form of backward iterations to accelerate the convergence toward samples from the distribution (see Propp and Wilson (1996)). More generally, the idea of backward dynamics is hidden in many constructions in Markov chain theory when the Markov chain is given by iterations of random operators as illustrated in Diaconis and Freedman (1999). In particular, they prove a general result concerning the convergence in distribution of these types of iterated Markov chains using the backward dynamics.
@@
Already instability issues appear in the full-batch regime, and a number of theoretical works have studied it under the heading edge of stability Cheng et al. (2020); Wu et al. (2024); Cai et al. (2024). Other works have also studied stability in the large batch regime after a batch size saturation takes place using the implicit conditioning ratio Lee et al. (2022); Agarwala and Pennington (2024). In this work we focus on the stochastic or small batch setting. Our findings do not really matter for the full-batch setting since backward and forward iterates then coincide. In the context of physics-informed neural networks, it has been observed that the gradient field is stiff, producing instabilities in learning trajectories Wang et al. (2020). To remedy this Li et al. (2023) propose a backward Euler scheme to stabilize training in this context. However the backward Euler method is an implicit Runge-Kutta method, which is different from using iteration backward.
@@
Curriculum learning Soviany et al. (2022) leverages the impact of sample order for generalization Mange (2019) by organizing training examples in a meaningful sequence, typically starting with simpler examples and gradually introducing harder ones, thereby optimizing the learning process and improving model performance. The backward optimization can be viewed as an automated form of curriculum, mitigating the forgetting of previous examples as new examples are added; a phenomenon which is also related to catastrophic forgetting and the stability gap in continual learning (see Lange et al. (2023) for instance).
@@ 2 A backward contraction principle
The contraction mapping principle, also called the Banach fixed point theorem, is a cornerstone result in mathematics and science, in particular for finding solutions of equations (Newton’s method) or of differential equations (Picard’s method of successive approximation). It is also behind Google’s PageRank algorithm Page et al. (1998). It concerns the existence of a fixed point of maps which are uniform contractions of a complete metric space [^1] Recall that a complete metric space $(\Omega,d)$ is a set $\Omega$, equipped with a distance metric $d(\theta_{1},\theta_{2})$ for which all Cauchy sequences (i.e., sequences $\theta_{n}$ such that $d(\theta_{n},\theta_{m})\rightarrow 0$ as $n,m\rightarrow 0$) converge to points in the space; typically $\mathbb{R}^{d}$ with the Euclidean distance is a complete metric space.. In detail, let $(\Omega,d)$ be a complete metric space and $T:\Omega\rightarrow\Omega$ be a continuous map. If there exists $0\leq k<1$ such that
@@
for any $\theta\in\Omega$ as $n\rightarrow\infty$. For example, in PageRank Page et al. (1998), one iterates the PageRank matrix from 50 to 100 times and the result is a very good approximation of the PageRank vector $\theta_{0}$. Another application is the standard proof of convergence for (full-batch) gradient descent around a minimum, which can be interpreted as an application of the Banach fixed point theorem as detailed in the next example:
@@
(Full-batch gradient descent convergence.) In this case the operator is $T(\theta)=\theta-h\nabla L(\theta)$ for a loss function $L$. The idea of the proof is to choose the learning rate $h$ small enough so that $T$ becomes a contraction. More precisely, around a minimum $\theta^{*}$ the operator $T$ can be approximated using a first-order expansion of the gradient around the minimum as $T(\theta)=\theta^{*}+(I-hH)(\theta-\theta^{*})$, where $H=\nabla^{2}L(\theta^{*})$. Now we have that
$\|T(\theta_{1})-T(\theta_{2})\|\leq\|I-hH\|_{\textrm{op}}\|\theta_{1}-\theta_{2}\|,$ (1)
where $\|I-hH\|_{\textrm{op}}$ is the operator norm of 1hH1-hH, that is the operator maximum eigenvalue: $\max_{i}|1-h\lambda_{i}|$ (here the $\lambda_{i}$’s are eigenvalues of $H$). Now it is easy to verify that $\|I-hH\|_{\textrm{op}}<1$ if and only if the learning rate is strictly smaller than $2/\lambda_{\textrm{max}}$ where $\lambda_{\textrm{max}}$ is the largest eigenvalue of $H$. Convergence for that setting follows from the Banach fixed point theorem.
The generalization of this convergence argument to Stochastic Gradient Descent (SGD) is actually problematic. The main reason is that now we do not have a single operator but a sequence of them $T_{1},T_{2},\dots$, each computing the loss gradient on a different batch of data. Even if each of the operators is a contraction it turns out that usual forward iterations will not converge to a point in general as illustrated in Example˜2.4. Theorem˜2.2, however, establishes the convergence of the backward iterations.
@@
Condition ˜2 expresses that the distance $d(\theta,T_{i}(\theta))$ is uniformly bounded by a constant $D$ for all $T_{i}^{\prime}s$ for a point $\theta$. Let us show that if this happens for a single point $\theta$, this happens for all points, provided we change the constant $D$. To see this, take another point $\tilde{\theta}$. We will compute another constant $\tilde{D}$ such that $d(\tilde{\theta},T_{i}(\tilde{\theta}))<\tilde{D}$. Namely, using the triangle inequality and the fact that $T_{i}$ are uniform contractions, we obtain that
$)\displaystyle d(\tilde{\theta},T_{i}(\tilde{\theta}$)) $\leq$ $)\displaystyle d(\tilde{\theta},\theta)+d(\theta,T_{i}(\theta))+d(T_{i}(\theta),T_{i}(\tilde{\theta}$))
@@
At this point, we can use the condition in ˜2 for the first term $d(\theta,T_{n+1}(\theta))<D$ and in conjunction with the contraction property for the subsequent terms in the sum, yielding:
@@
The contractivity assumption for the operators $T_{i}$’s is in general too restrictive for deep-learning settings. However, the proof of Theorem˜2.2 still works in a more generalized setting that encompasses the non-convexity of deep-learning loss-landscapes. The price to pay though is the introduction of the more abstract notion of a pseudo-metrics for which the argument in the proof of Theorem˜2.2 still works verbatim. We give the details of this more complicated approach in Appendix˜G.
As for the condition in ˜2, let us verify that it is satisfied for SGD in the case of an overparametrized model like a neural network. Namely, in this setting, we generally consider losses of the form $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}L_{i}(\theta)$, where $L_{i}$ is the loss computed on batch $B_{i}$ with $L_{i}(\theta)$ being non-negative. Because of the over-parameterization, there generally exists points $\theta^{*}$ where the full loss vanishes. This in turns implies that $L_{i}(\theta^{*})=0$ for all $i$. For SGD, the operators are $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$. Now it is easy to see that ˜2 is satisfied: $d(\theta^{*},T_{i}(\theta^{*}))=\|\theta^{*}-\theta^{*}-h\nabla L_{i}(\theta^{*})\|=0,$ since the global zero $\theta^{*}$ is a critical point for all the $L_{i}$’s.
Now here is a major point: The contraction mapping principle fails with the forward order. As the next example demonstrates, it is not enough for the operator $T_{i}$’s to be uniform contractions for the forward sequence of iterates to converge, while this is always true for the backward sequence because of Theorem˜2.2.
@@
(Forward iterations counter-example.) Here is an extreme example illustrating the convergence failure for the forward sequence of iterates (even when the maps are uniform contractions), while the backward sequence converges to a single point under the conditions of Theorem˜2.2. Consider the two constant maps $S(\theta)=x_{0}$, and $U(\theta)=y_{0}$, which are contractions with $k=0$. Assume that $x_{0}\neq y_{0}$, now also assume that each $T_{i}$ is either equal to $S$ or $U$ with equal probability to be selected. Then, independently of $\theta$, the forward iterates
@@
will jump between $x_{0}$ and $y_{0}$, according to whether $T_{n}$ is $S$ or $U$ at that particular $n$. Thus, no convergence to a single point is possible, although the sequence converges to a probability uniformly distributed in the two outcomes, since for each forward iterate either outcome has probability $1/2$. On the contrary, the backward trajectory